OpenStudy (anonymous):

4a=6-b 4a=5-b using substitution method

OpenStudy (anonymous):

\[4a=6-b \\4a=5-b\] ??

OpenStudy (anonymous):

doesn't seem very likely does it?

OpenStudy (anonymous):

how can \(4a\) be both \(5-b\) and at the same time \(6-b\) ?